Mass of CuSO4● XH2O (g) = 1.6 grams BEFORE HEATING

After Heating :

Mass of H2O(g) = 1.1g

Mass of CuSO4 (g) : 0.5


QUESTION:

Use the information to calculate the number of moles of water lost and the number of moles in the anhydrous compound



1
Expert's answer
2022-05-16T10:27:03-0400

"M(H_2O)=2x1+16=18 g\/mole"

"n(H_2O)=1.1\/18=0.061 mole"

"M(CuSO_4)=64+32+4x16=160 g\/mole"

"n(CuSO_4)=0.5\/160=0.0031 mole"

"n(H_2O)\/n(CuSO_4)=0.061\/0.0031=20"

So formula is "CuSO_4x20H_2O"